OpenStudy (anonymous):

I WILL FAN AND MEDAL!!!!!!!!!!!!!!!!!!!!!! Gary conducted a survey to find out the students' response to the cafeteria food. Fifty students think that there should be more healthy food served. This is five-sevenths of the total students surveyed. How many did Gary survey?

OpenStudy (anonymous):

50 students think that there should be healthy food served. let the total number of students that surveyed be X. 50 of the students is equal to 5/7 of the total students. there fore 50=5/7 *X X=50*7/5 X=70 students. there 70 students did the Gary survey.